OpenStudy (blackops2luvr):

Given the relation y = 3x^2 + 1, if the input is 4, what is the output? 13 48 49 145

OpenStudy (blackops2luvr):

@iGreen

OpenStudy (zale101):

y = 3x^2 + 1 f(4)=3(4)^2 + 1 f(4)= ?

OpenStudy (blackops2luvr):

....................................................

OpenStudy (igreen):

Input is x..so plug in 4 for x: y = 3x^2 + 1 y = 3(4)^2 + 1 Simplify exponent: y = 3(16) + 1 Multiply: y = 48 + 1 Can you add that?

OpenStudy (igreen):

Yep ^^

OpenStudy (blackops2luvr):

Which point could be removed in order to make the relation a function? {(–4, 3), (–5, 6), (1, 0), (–4, 5), (9, 5), (0, –7)} (–5, 6) (1, 0) (–4, 5) (9, 5)

OpenStudy (blackops2luvr):

@iGreen

OpenStudy (igreen):

Functions can't have two of the same x-value. So which one can you take out?

OpenStudy (blackops2luvr):

A

OpenStudy (blackops2luvr):

is it a? o.o @iGreen

OpenStudy (blackops2luvr):

the answer?

OpenStudy (igreen):

No, are there two x-values with -5? (1, 2) ^ x-value

OpenStudy (igreen):

There are two -4 x-values, so what's your answer gonna be?

OpenStudy (blackops2luvr):

c

OpenStudy (igreen):

Yep!

OpenStudy (igreen):

I g2g now..sorry. Bye!

OpenStudy (blackops2luvr):

grr i need to give yah a medal >.<

OpenStudy (igreen):

Eh, doesn't really matter :P